the portion of a line intercepted between the co-ordination axes is bisected at (3,-2) find the equation of the line.

Grade:12

1 Answers

Nehal Wani
13 years ago

 Let the line be x/a+y/b=1

The points of intersection of this line with the coordinate axes are (a,0) and (0,b)

Since (3,-2) bisects this line, it the the mid-point of the intercept.

Therefore, a/2=3 and b/2=-2

a=6 and b=-4

Thus the equation is : x/6-y/4=1 or 4x-6y=24 or 2x-3y=12